Please kindly solve this problem : Combine like radicals in 5sqrt2x-5sqrt32x^5+5sqrt-64x+3sqrt8x^3

Answer by Boreal(15235)

5sqrt2x-5sqrt32x^5+5sqrt64x+3sqrt8x^3; I edited sqrt (-64x) to sqrt (64x).
5sqrt(2x)
5sqrt 32x^5=4x^2 sqrt (2x)=20x^2 sqrt(2x); sqrt 32 is sqrt 16*sqrt2 and sqrt x^5 is x^2 sqrt x
5 sqrt(64x)=40sqrt(x)
3sqrt(8x^3)=6x(sqrt(2x), because sqrt 8=2 sqrt (2) and sqrt x^3=x sqrt(x)
Have 5 sqrt (2x)-20 x^2 sqrt(2x)+6x sqrt(2x)
and 40 sqrt(x)
The three terms are {-20x^2+6x+5){sqrt(2x} +40 sqrt (x)
Note, I could break down 64sqrt (x) to 4 sqrt(2)(sqrt(2x), but that would have two radicals.